### v1

# PhotoDoodle: Learning Artistic Image Editing from Few-Shot Pairwise Data

Abstract
We introduce PhotoDoodle, a novel image editing framework designed to facilitate photo doodling by enabling artists to overlay decorative elements onto photographs. Photo doodling is challenging because the inserted elements must appear seamlessly integrated with the background, requiring realistic blending, perspective alignment, and contextual coherence. Additionally, the background must be preserved without distortion, and the artist’s unique style must be captured efficiently from limited training data. These requirements are not addressed by previous methods that primarily focus on global style transfer or regional inpainting. The proposed method, PhotoDoodle, employs a two-stage training strategy. Initially, we train a general-purpose image editing model, OmniEditor, using large-scale data. Subsequently, we fine-tune this model with EditLoRA using a small, artist-curated dataset of before-and-after image pairs to capture distinct editing styles and techniques. To enhance consistency in the generated results, we introduce a positional encoding reuse mechanism. Additionally, we release a PhotoDoodle dataset featuring six high-quality styles. Extensive experiments demonstrate the advanced performance and robustness of our method in customized image editing, opening new possibilities for artistic creation. Code is released at https://github.com/showlab/PhotoDoodle.
{strip}[Uncaptioned image]
Figure 1: PhotoDoodle can mimic the styles and techniques of human artists in creating photo doodles, adding decorative elements to photos while maintaining perfect consistency between the pre- and post-edit states.
1 Introduction
The rise of diffusion models has started a new chapter for image creation and control. Using the pretrain-finetune approach, the community has achieved remarkable progress in customized image generation [25, 10, 12], with applications spanning identity preservation [18], artistic stylization [30, 45, 28, 2], and subject coherence [16, 25, 10, 12, 26, 14, 48]. However, while customized image generation thrives, a critical gap persists in customized image editing. This problem contrasts sharply with the growing need for precise image editing tools, as current methods mainly focus on creating new content rather than making smart edits based on context. This less explored area of custom image editing presents both technical challenges and chances to create new important applications.
Photo Doodling represents a long-recognized yet historically challenge in customized image editing. This task involves artists enhancing background photographs through the strategic integration of decorative elements and stylized modifications to achieve personalized or artistic effects. Typical techniques include localized stylization, decorative linework, novel object insertion, and ornamental augmentation of existing subjects. Artists and designers exhibit distinctive personal styles and strategic approaches during the photo doodling creation process, which holds significant value. However, the absence of automated tools forces manual workflows requiring hour-level time investments per image, severely constraining both productivity and the feasibility of generating large-scale paired training data.
Automating PhotoDoodle workflows is intriguing, yet it faces three interconnected challenges. First, achieving harmonious integration demands that generated decorations align not only spatially (e.g., matching perspective) but also semantically with the background, while accommodating non-photorealistic styles such as cartoon or flat illustration. Second, maintaining strict background consistency requires rigorously preventing unintended alterations to the original content, including color shifts or texture distortions. Third, to achieve efficient style acquisition, it is necessary to extract an artist’s unique editing patterns from a few photo-doodle pairs, which is technically challenging. These compounded challenges cause existing methods to be incompetent in addressing the problem comprehensively.
The limitations of prevailing image editing paradigms further exacerbate these difficulties. Global editing methods (e.g., Prompt-to-Prompt[11], InstructP2P[5]), while effective for consistent style transfer, inadvertently distort background content during localized modifications. Inpainting-based approaches, though capable of preserving non-mask regions through localized denoising, impose impractical demands for pixel-perfect user-defined masks, fundamentally conflicting with PhotoDoodle’s need for flexible creative workflows. To bridge this gap, we introduce an instruction-guided framework that replaces mask dependency with semantic-level control, enabling precise and style-conscious decorative generation while ensuring background consistency.
To address these challenges and bridge the gaps in existing solutions, we propose PhotoDoodle, a framework designed to learn artistic image editing from few-shot examples. Building upon a pre-trained Diffusion Transformer (DiT), our method conditions on background images to generate photo doodles through a two-stage training pipeline. First, we adapt a well-trained text-to-image model into a versatile image editor by fine-tuning it on large-scale editing datasets, thereby establishing the foundational editing model OmniEditor. Then, we train an EditLoRA module using artist-curated before-after pairwise data to learn nuanced editing styles, while introducing a Positional Encoding (PE) Cloning strategy to enforce strict background consistency through implicit feature alignment.
Our design is guided by two main insights. First, maintaining clean latent conditioning—ensuring that input image tokens are noise-free—is essential for keeping the output coherent with the background by minimizing divergence in the latent space. Second, implicit alignment through position encoding cloning enables the seamless integration of background context without adding extra trainable parameters, thus preserving the full generative capacity of the base T2I model. These insights allow our framework to strike a balance between artistic flexibility and strict consistency.
In summary, our contributions are threefold:
  • We propose a customizable image editing framework that learns and transfers artistic styles from few photo-doodle pairs.
  • By extending conditional mechanisms and implementing a PE cloning strategy for implicit alignment, we achieve high-quality and highly consistent image editing built upon Diffusion Transformers (DiTs), while utilizing EditLoRA to efficiently learn customized editing operations.
  • We propose the first dedicated photo-doodle dataset containing 300+ high-quality pairs across 6 artistic styles, establishing a benchmark for reproducible research.
2 Related work
2.1 Diffusion Model and Conditional Generation
Recent advances in diffusion models have significantly advanced the state of text-conditioned image synthesis, achieving remarkable equilibrium between generation diversity and visual fidelity. Pioneering works such as GLIDE [20], hierarchical text-to-image models [23], and photorealistic synthesis frameworks [27] have systematically addressed key challenges in cross-modal generation tasks. The emergence of Stable Diffusion [24], which implements a Latent Diffusion Model with text-conditioned UNet architecture, has established new benchmarks in text-to-image generation and inspired subsequent innovations including SDXL [22], GLiGEN [17], and Ranni [9]. To enhance domain-specific adaptation, parameter-efficient fine-tuning techniques like Low-Rank Adaptation (LoRA) [12] and DreamBooth [25] have demonstrated effective customization of pre-trained models. Concurrently, research efforts have focused on precise control over pictorial concepts through multi-concept customization [kumari2023multi], image prompt integration [39], and identity-preserving generation [36]. The introduction of ControlNet [43] further extended controllable generation capabilities to spatial constraints and depth information, with subsequent optimizations improving inference efficiency [46, 47]. Some work [35, 7, 32, 33] also focuses on the security issues of customized generation. Building upon these foundations, our work investigates the novel application of pre-trained diffusion transformers for generating artistic photo-doodles. Unlike previous approaches focusing on photorealism or explicit control modalities, we explore the model’s potential in capturing freehand artistic expression while maintaining structural coherence.
Figure 2: The overall architecture and training prodigim of photodoodle. The ominiEditor and EditLora all follow the lora training prodigm. We use a high rank lora for pre-training the OmniEditor on a large-scale dataset for general-purpose editing and text-following capabilities, and a low rank lora for fine-tuning EditLoRA on a small set of paired stylized images to capture individual artists’ specific styles and strategies for efficient customization. We encode the original image into a condition token and concatenate it with a noised latent token, controlling the generation outcome through MMAttention.
2.2 Text-guilded Image Editing
Recent advances in text-guided image editing have established this field as a critical research frontier in visual content manipulation, with current methodologies generally classified into three paradigms: global description-guided, local description-guided, and instruction-guided editing. Global description-guided approaches (e.g., Prompt2Prompt [11], Imagic [15], EditWorld [38] ), achieve fine-grained manipulation through cross-modal alignment between textual descriptions and image regions, yet demand meticulous text specification for target attributes. Local description-guided methods such as Blended Diffusion [3] and Imagen Editor [37] enable pixel-level control via explicit mask annotations and regional text prompts, though their practical application is constrained by the requirement for precise spatial specifications, particularly in complex editing scenarios like object removal. The emerging instruction-guided paradigm, exemplified by InstructPix2Pix [5] and HIVE [44], represents a paradigm shift through its natural language interface that accepts editing commands (e.g., ”change the scene to spring”). This approach eliminates the dependency on detailed textual descriptions or precise mask annotations, significantly enhancing user accessibility.
2.3 Image and video Doodles
Image and video doodling involve the creative process of adding hand-drawn elements or animations to static images or video content, blending aspects of graphic design, illustration, and animation to produce playful and engaging visual styles. In academic research, advanced techniques[40, 41, 4] have been developed to automate parts of this process. These methods enable users to generate intricate doodle animations from simple textual descriptions, sketches, or keyframes. By preserving artistic intent and reducing the time and expertise required, these models make multimedia content creation more accessible and efficient. Video doodling extends these capabilities to dynamic video sequences, allowing static doodle elements to be seamlessly integrated and animated in response to video motion and context. This innovation not only enhances interactivity but also introduces greater complexity and realism, making it a powerful tool for creative applications in entertainment, education, and storytelling.
3 Method
In this section, we begin by exploring the preliminaries on diffusion transformer as detailed in Section 3.1. Next, Section 3.2 outlines our three-stage system design. We then detail two key innovations—OmniEditor Pre-training (Section 3.3) and EditLoRA for style adaptation (Section 3.4). Finally, Section 3.5 describes how we built the PhotoDoodle dataset.
3.1 Preliminary
The Diffusion Transformer (DiT)[21] powers modern image generators like Stable Diffusion 3[8] and PixArt[6]. At its core, DiT uses a special transformer network that denoise the noisy image tokens step-by-step.
DiT operates on two categories of tokens: noisy image tokens $z\in\mathbb{R}^{N\times d}$ and text condition tokens $c_{T}\in\mathbb{R}^{M\times d}$, where $d$ denotes the embedding dimension, and $N$ and $M$ represent the numbers of image and text tokens, respectively. These tokens retain their shapes consistently as they propagate through multiple transformer layers.
In FLUX.1, each DiT block incorporates layer normalization, followed by Multi-Modal Attention (MMA) [8], which employs Rotary Position Embedding (RoPE) [34] to encode spatial information. For image tokens $z$, RoPE applies rotation matrices based on the token’s position $(i,j)$ in the 2D grid:
$z_{i,j}\rightarrow z_{i,j}\cdot R(i,j),$ (1)
where $R(i,j)$ represents the rotation matrix corresponding to position $(i,j)$. Similarly, text tokens $c_{T}$ are transformed with their positions designated as $(0,0)$.
The multi-modal attention mechanism projects these position-encoded tokens into query $Q$, key $K$, and value $V$ representations, enabling attention computation across all tokens:
$\text{MMA}([z;c_{T}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (2)
where $\mathit{Z}=[z;c_{T}]$ denotes the concatenation of image and text tokens. This approach ensures bidirectional attention between the tokens.
3.2 Overall Architecture
The overall architecture of PhotoDoodle is illustrated in Fig. 2. It comprises the following stages:
Pre-training OmniEditor. The OmniEditor is pre-trained on a large-scale image editing dataset to acquire general-purpose image editing capabilities and strong text-following abilities. This stage ensures the model’s capability in diverse editing tasks.
Fine-tuning EditLoRA. After pre-training, EditLoRA is fine-tuned on a small set of pairwise stylized images (20–50 pairs) to learn the specific editing styles and strategies of individual artists. This stage enables efficient customization for personalized editing needs.
Inference.: During inference, the input source image $I_{src}$ is encoded as condition tokens $c_{I}$ via VAE. We then randomly sample Gaussian noise as image tokens $z$, cloning the Position Encoding from condition tokens, and concatenate the tokens along the sequence dimension. Subsequently, we apply the flow matching method to predict the target velocity, iterating multiple steps to obtain the predicted image latent representation. Finally, the predicted image tokens are converted by the VAE decoder to achieve the final predicted image.
3.3 OmniEditor Pre-training
We denote the pre-edited image as the source image $I_{src}$ and the post-edited image as the target image $I_{tar}$. Previous works, such as SDEdit, model image editing as an adding-denoising problem, they often altering unintended areas. Others like InstructP2P[5] redesign core model parts, significantly degrading the capacity of the pretrained t2i models. Unlike them, our approach, PhotoDoodle, conceptualizes image editing as a conditional generation problem and minimizes the modification of pretrained text-to-image DiT.
Our model leverages the advanced capabilities of the DiT-based pretrained model, and extends it to function as an image editing tool. Both $I_{src}$ and $I_{tar}$ are encoded into their respective latent representations, $c_{I}$ and $z$, via a VAE. After applying position encoding cloning, the latent tokens are then concatenated along the sequence dimension to perform joint attention. The Multi-modal attention mechanisms are used to provide conditional information for the denoising of the doodle image.
$\text{MMA}([z;c_{i};c_{T}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (3)
where $\textit{Z}=[z;c_{i};c_{T}]$ denotes the concatenation of noised latent tokens, image condition tokens, and text tokens. Here, $c_{I}$ corresponds to $I_{src}$. This formulation enables bidirectional attention, letting the conditional branch and denoise branch interact on demand.
Position Encoding Cloning. Existing approaches to conditional image editing often struggle with pixel-level misalignment between input $I_{src}$ and edited output ($I_{tar}$) that undermine visual coherence. To address this fundamental challenge, we propose a novel Position Encoding(PE) Cloning strategy, motivated by the need for implicit spatial correspondence.
The PE Cloning is a simple yet powerful stragegy, it simply apply the position encoding calculated on $I_{src}$ on both $I_{src}$ and $I_{tar}$. The identical positional encoding serves as a strong hint so that the DiT is capable of learning correct corresponding easily. By enforcing identical positional encodings between the latent representations $c_{I}$ and $z$, our method establishes a pixel-perfect coordinate mapping that persists throughout the diffusion process. This geometric consistency ensures that every edit respects the original image’s spatial structure, eliminating ghosting artifacts and misalignments that plague conventional approaches.
Noise-free Conditioning Paradigm. A critical innovation lies in our noise-free conditioning paradigm. We preserve $c_{I}$ as a reference during the generation of $I_{tar}$. This design choice achieves two objectives through its operational duality. First, by maintaining $c_{I}$ in a noise-free state, we ensure the retention of high-frequency textures and fine structural details from the original image, thereby preventing degradation during iterative denoising. This preservation mechanism acts as a safeguard against the blurring artifacts commonly observed in conventional approaches. Second, the MM attention machanism is flexible enough to choose either to copy from the source or generate new content via instruction, making the model learns to manipulate only designated target regions.
Through the combined action of position encodings cloning and MMA mechanism, our framework achieves unprecedented precision in localized editing while maintaining global consistency, a balance previously unattainable in conditional image generation tasks.
Conditional flow matching loss. The conditional flow matching loss function is following SD3 [8], which is defined as follows:
$L_{CFM}=E_{t,p_{t}(z|\epsilon),p(\epsilon)}\left[\left\|v_{\Theta}(z,t,c_{I},c_{T})-u_{t}(z|\epsilon)\right\|^{2}\right]$ (4)
Where $v_{\Theta}(z,t,c_{I},c_{T})$ represents the velocity field parameterized by the neural network’s weights, $t$ is timestep, $c_{I}$ and $c_{T}$ are image condition tokens extracted from source image $I_{src}$ and text tokens. $u_{t}(z|\epsilon)$ is the conditional vector field generated by the model to map the probabilistic path between the noise and true data distributions, and $E$ denotes the expectation, involving integration or summation over time $t$, conditional $z$, and noise $\epsilon$.
3.4 EditLoRA
LoRA [12] enhances model adaptation by freezing the pre-trained model weights $W_{0}$ and inserting trainable rank decomposition matrices $A$ and $B$ into each layer of the model. These matrices, $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$, where $r\ll\min(d,k)$, are used to fit the residual weights adaptively. The forward computation integrates these modifications as follows:
$y^{\prime}=y+\Delta y=W_{0}x+BAx$ (5)
where $y\in\mathbb{R}^{d}$ is the output and $x\in\mathbb{R}^{k}$ denotes the input. $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$ with $r\ll\min(d,k)$. Normally, matrix $B$ is initialized with zero.
To learn an individual artist’s editing style and effectively transfer it from a small number of before-and-after image pairs, we introduce EditLoRA. Inspired by recent low-rank adaptation (LoRA) techniques [31, 29], EditLoRA fine-tunes only a small set of trainable parameters, significantly reducing the risk of overfitting while preserving most of the pretrained model’s expressive power. In our work, the general-purpose OmniEditor is trained on a large-scale paired dataset with a higher-rank lora. The EditLoRAs are lower-rank loras that specifically focus on mimicking the style and strategies of single artists in creating photo doodles. EditLoRA’s training set consists of before-and-after pairwise data and corresponding text instruction, which differ from the conventional text-image pairs required for learning image generation models.
The EditLora steers the behaviour of the OmniEditor to the specified artist’s style. When a new image $I_{src}$ is provided, along with the textual instructions, the model generates $I_{tar}$ that reflects both the previously learned editing capabilities and the distinctive stylistic effects from the artist.
Figure 3: The generated results of PhotoDoodle. PhotoDoodle can mimic the manner and style of artists creating photo doodles, enabling instruction-driven high-quality image editing.
Figure 4: Compared to baseline methods, PhotoDoodle demonstrates superior instruction following, image consistency, and editing effectiveness.
3.5 PhotoDoodle
In collaboration with professional artists and designers, we created the first PhotoDoodle dataset. We introduce the dataset containing six high-quality styles and over 300 photo doodle samples. The six styles include cartoon monster, hand-drawn outline, 3D effects, flowing color blocks, flat illustrator, and cloud sketch. Each sample in our dataset consists of a pre-edit photo (e.g., a real-world scene or portrait) and a post-edit photo doodle showing unique modifications by the artist, such as localized stylization, decorative lines, newly added objects, or modifications to existing elements. For each example, we store the raw input image $I_{src}$ and the doodled version $I_{tar}$, along with textual instructions.
4 Experiment
4.1 Experiment Setting
Setup. During the OmniEditor pre-training stage, we take the parameters of Flux.1 dev model as the initialization of the DiT architecture, and trained it with the SeedEdit dataset. Images were resized to 768x512. We trained a LoRA rank of 256, a batch size of 128, and a learning rate of $1\times 10^{-4}$, on 8 H100 GPUs for 330,000 steps. After merge the lora into the base DiT model, we acquired the OmniEditor model for further usage. In the EditLoRA training phase, we fine-tuned the merged model on a paired photo doodle dataset (50 pairs) using a single GPU for 10,000 steps, with a LoRA rank of 128, batch size of 2, and a learning rate of $1\times 10^{-4}$.
Baseline Methods. The baselines compared in this paper include InstructP2P[5], MagicBrush[42], and SDEdit[19] based on Flux. For a fair comparison, tests were conducted in both general image editing and customized editing scenarios. For the general image editing tests, OmniEditor was evaluated against the aforementioned baselines. In the customized editing scenario, we trained Flux LoRA using doodle results created by professional artists and used it alongside SDEdit as a baseline. For InstructP2P and MagicBrush, the attention layers were also fine-tuned with the same doodle dataset. Finally, all the trained LoRA models were compared with the proposed EditLoRA model.
Benchmarks. As with previous methods, we tested the performance of the proposed OmniEditor on the HQ-Edit benchmark[13]. For the customized generation tasks, this paper introduced a new benchmark collected from the Pexels website, consisting of 50 high-quality photographs of portraits, animals, and architecture.
4.2 Generation Results
Fig. 3 displays the image editing results of PhotoDoodle, which excels in text following due to training on a large dataset of before-after pairs. The generated doodles blend well with the original images. When trained on a limited dataset of artist-paired data using EditLoRA, PhotoDoodle consistently produces artistic doodles while preserving image consistency and avoiding unwanted changes. Notably, our method maintains stable performance and a high success rate, making it suitable for production use and reducing the need for selective sampling.
Table 1: Comparison Results in General Image Editing Tasks. The best results are denoted as Bold.
Methods $CLIP\ Score$ $GPT\ Score$ $CLIP_{img}$
Instruct-Pix2Pix 0.237 38.201 0.806
Magic Brush 0.234 36.555 0.811
SDEdit(FLUX) 0.230 34.329 0.704
Ours 0.261 51.159 0.871
4.2.1 Qualitative Evaluation
In this section, we present the results of the qualitative analysis. As illustrated in Fig. 4, OmniEdito demonstrates superior consistency and minimizes unintended alterations in general image editing tasks compared to state-of-the-art (SOTA) methods. This performance is attributed to the use of high-quality datasets and a thoughtfully designed model architecture. For custom image editing tasks, our method significantly surpasses baseline methods, as evidenced by the high quality of generated outputs and the strong alignment of the editing style with the original artistic intent, while avoiding undesired modifications.
Table 2: Comparison Results in Customized Image Editing Tasks. The best results are denoted as Bold.
Methods $CLIP\ Score$ $GPT\ Score$ $CLIP_{img}$
Instruct-Pix2Pix 0.249 36.359 0.832
Magic Brush 0.247 38.478 0.885
SDEdit(FLUX) 0.209 21.793 0.624
Ours 0.279 63.207 0.854
Figure 5: Ablation study results. Without OmniEditor pre-training, training EditLoRA directly reduces the harmony between sketches and photos and weakens text-following capabilities; removing Position Encoding Cloning decreases output consistency and introduces unwanted background changes; using only pre-trained OmniEditor without EditLoRA significantly reduces the degree of stylization.
4.2.2 Quantitative Evaluation
In this section, we present the quantitative analysis results. Following InstructP2P[5], we compute the $CLIP\ Score$ and $CLIP_{img}$ metrics for both tasks. Furthermore, as proposed in HQ-Edit [13], we employ GPT4-o to evaluate the alignment between text instructions and editing outputs. As shown in Table 1, our method outperforms all baselines across all metrics in general image editing tasks, achieving the highest $CLIP\ Score$, $GPT\ Score$, and $CLIP_{image}$ Score. In custom image editing tasks, while some models fail to produce meaningful edits, which leads to high $CLIP_{image}$ scores, our method still holds a clear advantage over the baselines. This is evident in the substantial improvements in $GPTScore$ and $CLIPScore$, both of which evaluate the consistency and quality of the generated content in relation to the artist’s original work.
4.3 Ablation Study
To demonstrate the effectiveness of the key strategies and modules proposed in this paper, we conducted detailed ablation experiments. We evaluated OmniEditor Pre-training, Position Encoding Cloning, and EditLoRA. As shown in Fig. 5: Without OmniEditor Pre-training, directly training EditLoRA leads to reduced harmony between the generated sketches and photos, along with weaker text-following capabilities in the output. Removing Position Encoding Cloning results in decreased consistency in the generated outputs, with unwanted changes occurring in the background. When EditLoRA is not used, and only the pre-trained OmniEditor is employed for generation, the degree of stylization in the results is significantly reduced.
4.4 User Study
To further demonstrate the superiority of our proposed method, we conducted a user study with 30 participants via online questionnaires. We evaluatedd user preferences in both general and customized image editing scenarios. Participants were presented with PhotoDoodle’s outputs alongside baseline methods, and asked to evaluate which results they preferred based on three criteria: 1) Overall preference, 2) Instruction following, and 3) Consistency between the edited images and the original images. During the study, participants viewed the original unedited images, the edit instructions, and reference images edited by models. They were then asked to decide whether PhotoDoodle (Option A) or a baseline method (Option B) performed better, or if they were about equally effective. The results of this user study are collected in Fig. 6, where we reported the percentage scores of each criterion, highlighting our method’s effectiveness in aligning closely with artistic intentions and maintaining high consistency in edits without introducing unwanted changes.
Figure 6: User study results. The scores demonstrate the percentage of users who prefer ours over others under three evaluation metrics. PhotoDoodle outweighs all other baselines in user study.
5 Limitation and Future Work
One limitation of PhotoDoodle is its dependence on the collection of dozens of paired datasets (pre-edit and post-edit images) and the need for thousands of training steps using LoRA. This data collection process can be challenging, as paired images are not always readily accessible. In the future, we will attempt to learn doodling strategies from single image pairs using an Encoder structure.
6 Conclusion
In this paper, we present PhotoDoodle, a diffusion-based framework for artistic image editing that learns unique artistic styles from minimal paired examples. By combining large-scale pretraining of the OmniEditor with efficient EditLoRA fine-tuning, PhotoDoodle enables precise decorative generation while maintaining background integrity through positional encoding cloning. Key innovations, including a noise-free conditioning paradigm and parameter-efficient style adaptation requiring only 50 training pairs, significantly reduce computational barriers. We also contribute a new dataset with six artistic styles and 300+ curated samples, establishing a benchmark for reproducible research. Extensive experiments demonstrate superior performance in style replication and background harmony, outperforming existing methods in both generic and customized editing scenarios.